OpenStudy (studybro):

If two angles of a triangle have the same measure and the third angle is 72 degrees greater than the measure of the other two. Find the measure of each angle. x=? x+72=?

OpenStudy (bradely):

In triangle total angle 180 2x+x+72=180 3x=180-72 x=36 36,36,108 Source: http://www.mathskey.com/question2answer/

OpenStudy (studybro):

thank you thank you thank you!!!!!!(: